x+7y=-37
2x-5y=21
how do i solve this by using the method of elimination. If Step by step process would be shown, it would be greatly appreciated.
Thank you :)

Found 2 solutions by Alan3354, enow:
Answer by Alan3354(69443)

x+7y=-37
2x-5y=21
========
Multiply the 1st eqn by 2, then subtract.
That eliminates x

Answer by enow(2)

I'll suggest you to multiply the equation,x+7y=-37 with 2 so that you can get the same coefficient of x as the other equation. then, work can be easier for you!
x2) x+7y=-37 ; 2x-5y=21---Eqn. 2
2x+14y=-74---Eqn. 1
By method of elimination, Eqn. 1 - Eqn. 2:
(2x+14y)-(2x-5y)=-74-21
19y=-95
therefore, y=-5.
Substitute y=-5 into any of the equations, then you will get x=-2 as your answer.
Hope I helped!